# Work performed during an internship at Baidu

### v1

@@ 4 Experiments
To validate the effectiveness of our work, we chose three modalities and conducted a comprehensive evaluation of our approach on five datasets. Experimental results demonstrate that SegICL exhibits outstanding in-context learning capabilities and delivers competitive results across multiple datasets.
@@ 4.1 Dataset
71 publicly available datasets is utilized as the training set, referred to as the in-distribution datasets (details can be found in the supplementary materials).
Additionally, we reserved the fundus segmentation task to evaluate SegICL’s learning capabilities for OOD tasks. Furthermore, for each modality, we retained some datasets to assess the performance of SegICL on OOD tasks. Regarding the training set, we preserved a portion as the test set for evaluating the model’s segmentation performance on in-distribution data.
The evaluation perform on three fundus datasets, along with one CT dataset and one MR dataset. Additionally, we applied the approach described in [44] to uniformly preprocess the experimental datasets.
REFUGE2[14]. This dataset contains 2000 color fundus images with annotations of glaucoma classification, optic disc/cup segmentation, as well as fovea localization. We conducted evaluations on the optic disc/cup segmentation task.
PALM[13]. This dataset consists of 1200 images with labels for pathologic myopia, including manual annotations for the optic disc, fovea position. We chose evaluations on the optic disc segmentation task.
IDRiD[27]. This dataset includes disease severity information for diabetic retinopathy and diabetic macular edema in each image. There are five segmentation tasks, and we chose optic disc segmentation for the evaluation experiment.
CHAOS[18]. This dataset includs abdomen MRI data, which sourced from ISBI 2019 Combined Healthy Abdominal Organ Segmentation Challenge (Task 5). We focus on four segmentation classes:liver, left kidney, right kidney, spleen.
BTCV[21]. This dataset includes a total of 50 abdominal CT scans, obtained from patients with metastatic liver cancer or postoperative abdominal hernia. It involves 13 segmentation categories, and we adhere to the aforementioned evaluation metrics, focusing specifically on four segmentation classes: liver, left kidney, right kidney and spleen.
@@ 4.2 Implementation details
SegICL utilizes QWen-7B as the multi-modal encoder and ControlNet as the image decoder. A 4-layer MLP (Multi-Layer Perceptron) is used as both the condition encoder and the teacher module, with shared parameters between them. After the completion of training in the first stage, the parameters of condition encoder and the teacher module are frozen.
In the first stage, the image decoder is trained using 2M image mask data, and after training, all parameters of the image decoder are frozen. In the second stage, a two-stage instruction dataset is employed to train the multi-modal encoder, covering a total of 71 datasets, including 5 modalities, with a total of 200K instruction data. LoRA[16] fine-tuning is applied to the multi-modal encoder, while Full-parameter finetuning is employed for the projector part.
The Dice[10] score is used as the evaluation metric for the model. For the calculation of few-shot performance, we randomly sample data from the training set of the dataset under evaluation as example data. The final result is obtained by averaging the pixel-wise results of 5 independently randomly sampled predictions, followed by post-processing through binarization of the output masks.
@@ 4.3 Performance of the OOD modality
This section primarily analyzes modalities and their segmentation tasks beyond the distribution of the training dataset. The fundus image data did not utilized in the training dataset for SegICL.
Through experiments, it is found that multimodal large models naturally understand fundus images as shown in Fig. 4, and they can know the approximate shape and position of the optic disc. Therefore, different base models may lead to different performance changes, which is also a future direction of our work. Notably, our attained generalization capability is achieved without additional training and demands only a handful of relevant examples for substantial performance improvement.
We evaluated the effectiveness of SegICL on three datasets.(See Fig. 5) On the REFUGE2 dataset, employing SegICL-0 segmentation by directly using text prompts for optic disc and cup segmentation tasks resulted in scores of 0.447 and 0.294, respectively. This indicates that OOD tasks pose a challenging endeavor, requiring models to perform segmentation tasks effectively even in the absence of examples.
For the optic and disc segmentation task, introducing SegICL-1 lead to a performance improvement of 0.245, using SegICL-2 further increased performance by 0.059, and providing SegICL-2 resulted in an additional performance improvement of 0.088. It can be observed that with the addition of examples, SegICL can effectively capture the feature correlations between contexts, thereby transitioning OOD tasks from being undoable (poor performance) to achievable albeit with suboptimal performance.
Regarding the optic cup task, introducing SegICL-1 lead to performance improvements of 0.177, using SegICL-2 increased performance by 0.122, and providing SegICL-3 lead to a performance improvement of 0.079. On the PALM dataset, a similar trend is observed, with the Dice score increasing from 0.385 in SegICL-0 to 0.816 in SegICL-3, indicating that as the number of examples increases, the model gradually adapts to new segmentation tasks. On the IDRiD dataset, as the number of examples increased, the performance also gradually improved, with increments of 0.218, 0.093, and 0.084, respectively.
@@
Based on the above experiments, it can be concluded that SegICL possesses powerful in-context learning capabilities. When confronted with a new task without prior training, the segmentation performance is initially low. However, through example-based teaching, the model can grasp the segmentation approach for the task (See Fig. 6). With more examples provided for teaching, the model’s segmentation performance on that task shows a linear growth. Furthermore, the observed performance enhancement exhibits a positive correlation with the number of examples provided (See Fig. 6 SegICL-0 to SegICL-3).
Figure 6: The qualitative analysis diagram of SegICL on fundus segmentation tasks. As more examples are provided, the shape and position of the masks increasingly approximate the ground truth.
@@ 4.4 Performance of the OOD dataset
The generalization performance of SegICL is evaluated using an abdominal MR image dataset that did not appear in the training dataset. However, despite the presence of segmentation task data similar to the training dataset in this OOD dataset, they are not sampled from the same data distribution.
@@
PANet[37] 1-shot 50.40 40.58 32.19 30.99 38.53
SE-Net[31] 1-shot 29.02 47.30 47.96 45.78 42.51
ALPNet[25] 1-shot 62.35 61.32 60.81 58.83 63.17
RPNet[34] 1-shot 73.51 69.85 70.00 70.48 79.26
GCN-DE[32] 1-shot 49.47 60.63 83.03 76.07 67.03
VQNet[17] 1-shot 81.72 79.08 68.94 60.03 72.44
CRAPNet[11] 1-shot 76.46 74.32 81.95 86.42 79.79
SSL-ALPNet[25] 1-shot 76.10 72.18 85.18 81.92 78.84
SSL-VQNet[17] 1-shot 79.92 77.21 91.56 89.54 84.56
@@
Table 1: Performance comparison of SegICL and SOTA few-shot models on MRI dataset, SSL represents the utilization of self-supervised learning to enhance model performance, while SegICL-x stands for the number of samples for in-context learning.
In Tab. 1, we compared the performance of SegICL with SOTA few-shot methods. All few-shot methods are based on 1-way 1-shot settings. As shown in the table, our approach achieved an average Dice score higher than VQNet, by 1.52% in the SegICL-0 setting. Specifically, our SegICL-0 outperformed VQNet by 13.26% and 20.44% in Kidney segmentation. Compared to incorporating Self-Supervised Learning (SSL), SegICL, with the introduction of In-Context Learning (ICL), demonstrated further improvement in performance. Its average Dice score surpassed the strongest few-shot method, VQNets, by 0.57%. Moreover, it reached SOTA levels in spleen, left kidney, and right kidney sub-segmentation tasks. It is important to note that our method achieved inference in a completely train-free manner, avoiding the need to retrain the network with an SSL module.
@@
Additionally, the table reflects the necessity of In-Context Learning for performance enhancement on OOD tasks. It is evident that, for the four sub-classification tasks, performance improved by 8.56%, 15.97%, 9.98%, and 9.48%, respectively, resulting in an average Dice improvement of 11.18%. Thus, SegICL demonstrates robust generalization performance, achieving competitive results on OOD tasks with the incorporation of In-Context Learning.
SegICL demonstrates a certain level of generalization ability for OOD datasets and segmentation tasks. Moreover, as the number of example data increases, SegICL exhibits a capacity for learning new knowledge, with performance improving proportionally to the increase in example data.
@@
Figure 8: Performance comparison on CT dataset. The performance of SegICL is comparable to that of models tailored specifically for this dataset.
@@ 4.5 Performance of the In-distribution datasets
In this section, we conducted performance evaluation of SegICL using the test set from the dataset that is previously used in the training set. As shown in Fig. 8, SegICL exhibits comparable performance to models trained specifically on dedicated datasets. We used nnU-Net, trained on the BTCV dataset through supervised learning, as the upper performance limit. In SegICL-0 inference, SegICL’s performance differs from UNet by only 0.4%. After employing in-context learning with example data selected from the training set, our method’s average performance surpasses UNet by 0.5%, demonstrating the learning capability of SegICL.
Simultaneously, the table indicates that in this experimental setup, the learning capacity of SegICL is noteworthy. Through in-context learning, the model is able to capture contextual features of the segmentation task, leading to improved segmentation performance.
SegICL falls short of surpassing proprietary models trained specifically for dedicated datasets due to its training dataset encompassing multiple modalities from various datasets. As a result, its performance on a single dataset can only be competitive rather than achieving SOTA status. The primary purpose behind introducing SegICL is to propose a method capable of addressing OOD tasks through contextual learning with a limited number of examples. Achieving the utmost optimization in performance is a consideration for future work.
@@ 4.6 Qualitative results
For a more in-depth exploration of SegICL’s segmentation performance, we visualize four typical CT images featuring the liver, spleen, kidneys, and pancreas in Fig. 9. The results reveal that SegICL excels in segmenting CT slices with larger and more regular-shaped target volumes, such as the upper part displaying the liver and kidneys. However, its performance diminishes on CT slices with smaller and elongated target shapes. The segmentation edge localization capability is inadequate for all CT slices.
Qualitative analysis unveils two potential issues. Firstly, the limited modeling capacity of the multimodal encoder for the state vector. While larger models could enhance fitting, we opted for the 7B language model considering computation costs and model effectiveness. Secondly, the use of diffusion to generate segmentation masks inherits issues from the diffusion model, leading to a slight deficiency in generating fine-grained segmentation details. Nonetheless, existing medical segmentation works based on diffusion[41] have demonstrated its significant improvement in segmentation performance, maybe this issue can be solve. In summary, SegICL is a successful universal medical segmentation framework with robust context learning capabilities. The experiments confirm its efficacy. Future research will explore solutions to the aforementioned issues.
Figure 9: Qualitative analysis results display: Segmentation results for different organs in CT imaging. Demonstrate the strengths and limitation of SegICL.
5 Discussion
This paper has several limitations: Firstly, due to our use of a pre-trained language model, the token length of inputs and outputs is limited by the settings of the base model, which prevents us from exploring the performance upper limit of SegICL-x. Secondly, there are issues with poor edge processing for the generated masks, even with post-processing, precise edge segmentation cannot be achieved. Therefore, it is necessary to explore new technologies to increase segmentation accuracy by increasing the token length and refining the mask edge.
Future research can be focused on the following three aspects: Firstly, training the model using larger models and datasets. Secondly, enhancing the quality of image generation and segmentation performance. From the above analysis ( Sec. 4.6), it can be inferred that the segmentation performance of the model is related to the quality of image generation. Therefore, improving the details of generated images is also crucial for enhancing the performance of the model. Thirdly, accelerating the segmentation speed of the model. Due to the introduction of large language models and Diffusion models, the segmentation speed is relatively slow, so it is necessary to study how to improve the speed of the model.
6 Conclusion
In this study, we introduce SegICL, a pioneering approach to image segmentation leveraging text-guided segmentation and In-Context Learning (ICL) that effectively addresses the limitations of universal segmentation models on OOD task with a training-free manner. Our work paves the way for more intuitive, efficient, and effective segmentation models that can adapt to the rapidly evolving landscape of medical imaging.

### v4

@@ 4.1 Dataset
The training set includes 71 publicly available datasets. Additionally, we reserved the optic fundus segmentation task to evaluate SegICL’s learning capabilities for OOD tasks. Furthermore, for each modality, we retained some datasets to assess the performance of SegICL on OOD tasks. Additionally, we applied the approach described in a27 to uniformly preprocess the experimental datasets.
The evaluation perform on three optic fundus datasets, along with one CT dataset and one MRI dataset. REFUGE2a32 . This dataset contains 2000 color optic fundus images with annotations of glaucoma classification, optic disc/cup segmentation, as well as fovea localization. We conducted evaluations on the optic disc/cup segmentation task. PALMa33 . This dataset consists of 1200 images with labels for pathologic myopia, including manual annotations for the optic disc, fovea position. We chose evaluations on the optic disc segmentation task. IDRiDa34 . This dataset includes disease severity information for diabetic retinopathy and diabetic macular edema in each image. There are five segmentation tasks, and we chose optic disc segmentation for the evaluation experiment. CHAOSa35 . This dataset includs abdomen MRII data, which sourced from ISBI 2019 Combined Healthy Abdominal Organ Segmentation Challenge (Task 5). We focus on four segmentation classes:liver, left kidney, right kidney, spleen. BTCVa36 . This dataset includes a total of 50 abdominal CT scans. We adhere to the aforementioned evaluation metrics, focusing specifically on four segmentation classes: liver, left kidney, right kidney and spleen.
@@ 4.2 Implementation details
Model training is divided into two stages (see appendix). LoRAa44 is applied to the multimodal encoder. The Dicea37 score is used as the evaluation metric for the model. For the calculation of few-shot performance, we randomly sample data from the training set of the dataset under evaluation as multimodal prompt. The final result is obtained by averaging the pixel-wise results of 5 independently randomly sampled predictions, followed by post-processing through binarization of the output masks.
@@ 4.3 Performance of the OOD modality
This section primarily analyzes modalities and their segmentation tasks beyond the distribution of the training dataset. The optic fundus image data did not utilized in the training dataset for SegICL. Pre-trained models understand optic fundus images as shown in Fig.4, and the approximate shape and position of the optic disc can be identified. We evaluated the effectiveness of the proposed framework on three datasets(See Fig.5). On the REFUGE2 dataset, employing SegICL-0 by directly using text prompts for optic disc and cup segmentation tasks resulted in scores of 0.447 and 0.294, respectively. This indicates that the prior knowledge of large models can provide some help for the segmentation of the model.
Figure 6: The qualitative analysis diagram of SegICL on optic fundus segmentation tasks. As more examples are provided, the shape and position of the masks increasingly approximate the GT.
For the optic disc and cup segmentation task, introducing SegICL-0 to SegICL-3 brings continuous performance growth. It can be observed that with the addition of shots, SegICL can effectively capture the feature correlations between contexts, thereby transitioning OOD tasks from being undoable to achievable albeit with suboptimal performance. Based on the above experiments, it can be concluded that SegICL possesses powerful in-context learning capabilities. When confronted with a new task without prior training, the segmentation performance is initially low. However, through example-based teaching, the model can grasp the segmentation approach for the task (See Fig.6). With more shots provided, the model’s segmentation performance on that task shows a linear growth.
@@ 4.4 Performance of the OOD dataset
The generalization performance of SegICL is evaluated using an abdominal MRI image dataset that did not appear in the training dataset. However, despite the presence of segmentation task data similar to the training dataset in this OOD dataset, they are not sampled from the same data distribution.
@@
PANeta38 1-shot 50.40 40.58 32.19 30.99 38.53
SE-Neta39 1-shot 29.02 47.30 47.96 45.78 42.51
ALPNeta7 1-shot 62.35 61.32 60.81 58.83 63.17
RPNeta18 1-shot 73.51 69.85 70.00 70.48 79.26
GCN-DEa41 1-shot 49.47 60.63 83.03 76.07 67.03
VQNeta17 1-shot 81.72 79.08 68.94 60.03 72.44
CRAPNeta51 1-shot 76.46 74.32 81.95 86.42 79.79
SSL-ALPNeta7 1-shot 76.10 72.18 85.18 81.92 78.84
SSL-VQNeta17 1-shot 79.92 77.21 91.56 89.54 84.56
@@
Table 1: Performance comparison of SegICL and SOTA few-shot models on MRII dataset, SSL represents the utilization of self-supervised learning to enhance model performance, while SegICL-x stands for the number of samples for in-context learning.
@@
In Tab.1, we compared the performance of SegICL with SOTA few-shot methods. All few-shot methods are based on 1-shot settings. As shown in the table, our approach achieved an average Dice score higher than VQNet, by 1.52% in the SegICL-0 setting. Specifically, our SegICL-0 outperformed VQNet by 13.26% and 20.44% in Kidney segmentation. Compared to incorporating Self-Supervised Learning (SSL), SegICL, with the introduction of In-Context Learning (ICL), demonstrated further improvement in performance. Its average Dice score surpassed the strongest few-shot method, VQNets, by 0.57%. Moreover, it reached SOTA levels in spleen, left kidney, and right kidney sub-segmentation tasks. It is important to note that our method achieved inference in a completely train-free manner, avoiding the need to retrain the network with an SSL module. Additionally, the table reflects the necessity of In-Context Learning for performance enhancement on OOD tasks. It is evident that, for the four sub-classification tasks, performance improved by 8.56%, 15.97%, 9.98%, and 9.48%, respectively, resulting in an average Dice improvement of 11.18%. Thus, SegICL demonstrates robust generalization performance, achieving competitive results on OOD tasks. SegICL demonstrates a certain level of generalization ability for OOD datasets and segmentation tasks.
@@ 4.5 Performance of the In-distribution datasets
Figure 8: Performance comparison on CT dataset. The performance of SegICL is comparable to those model specifically customized for this dataset.
In this section, we conducted performance evaluation of SegICL using the test set from the dataset that is previously used in the training set. As shown in Fig.8, SegICL exhibits comparable performance to models trained specifically on dedicated datasets. We used nnU-Net, trained on the BTCV dataset through supervised learning, as the performance benchmark. In SegICL-0 inference, SegICL’s performance differs from UNet by only 0.4%. After employing in-context learning with example data selected from the training set, our method’s average performance surpasses UNet by 0.5%, demonstrating the learning capability of SegICL.
Simultaneously, the table indicates that in this experimental setup, the learning capacity of SegICL is noteworthy. Through in-context learning, the model is able to capture contextual features of the segmentation task, leading to improved segmentation performance. SegICL falls short of surpassing proprietary models trained specifically for dedicated datasets due to its training dataset encompassing multiple modalities from various datasets. As a result, its performance on a single dataset can only be competitive rather than achieving SOTA status. The primary purpose behind introducing SegICL is to propose a method capable of addressing OOD tasks through ICL with a limited scale.
@@ 4.6 Qualitative results
For a more in-depth exploration of SegICL’s segmentation performance, we visualize four typical CT images featuring the liver, spleen, kidneys, and pancreas in Fig.9. The results reveal that SegICL excels in segmenting CT slices with larger and more regular-shaped target volumes, such as the upper part displaying the liver and kidneys. However, its performance diminishes with smaller and elongated target shapes. Qualitative analysis unveils two potential issues. First, the limited modeling capacity of the multimodal encoder for the state vector. Alternatively, increasing the number of examples can lead to better segmentation results. Secondly, the use of diffusion to generate segmentation masks inherits issues from the diffusion model, leading to a slight deficiency in generating fine-grained segmentation details. Nonetheless, existing segmentation works based on diffusiona52 have demonstrated its significant improvement in segmentation performance, maybe this issue can be solve.
Figure 9: Qualitative analysis results display: Segmentation results for different organs in CT image. Demonstrate the strengths and limitation of SegICL.
5 Discussion and Broader Impact
Discussion: First, due to computing resource constraint, the token length of inputs and outputs is limited, which prevents us from exploring the performance upper limit of SegICL-x. Second, there’s still room for improvement in contour processing for the generated masks, even with post-processing, precise contour segmentation cannot be achieved. Future research can be focused on the following three aspects: Firstly, training the model using larger models and datasets, transitioning from Few-Shot to Many-Shot ICL. Secondly, improving the details of generated images is also crucial for enhancing the performance of the model. Thirdly, accelerating the segmentation speed of the model. Due to the introduction of large language models and diffusion models, the segmentation speed is relatively slow, so it is necessary to study how to improve the speed of the model.
Broader Impact: In this paper, we introduce SegICL, a pioneering approach to image segmentation leveraging text-guided segmentation and ICL that effectively addresses the limitations of universal segmentation models on OOD task with a training-free manner. Our work redefine the paradigms for segmentation tasks in the era of larger models and paves the way for more intuitive, efficient, and effective segmentation models that can adapt to the rapidly evolving landscape of medical image.